# Two-Player Zero-Sum Games with Bandit Feedback

### v1

@@ Abstract
We study a two-player zero-sum game (TPZSG) in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback. We propose and analyze two algorithms: ETC-TPZSG, which directly applies ETC to the TPZSG setting and ETC-TPZSG-AE, which improves upon it by incorporating an action pair elimination (AE) strategy that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair. Our objective is to demonstrate the applicability of ETC in a TPZSG setting by focusing on learning pure strategy Nash Equilibrium. A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret for both algorithms, has received limited attention in the literature on zero-sum games. Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC-TPZSG and $O(\frac{\log(T\Delta^{2})}{\Delta})$ for ETC-TPZSG-AE, where $\Delta$ denotes the suboptimality gap. Therefore, our results indicate that ETC-based algorithms perform effectively in adversarial game settings, achieving regret bounds comparable to existing methods while providing insights through instance-dependent analysis.
@@ 1 Introduction
This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in an adversarial setting with strictly opposing objectives, as first analyzed in v1928theorie ; vonneumanandmorgenstern . These games play a foundational role in game theory and online learning, capturing adversarial interactions in various applications such as economic competition and adversarial machine learning. In such settings, players repeatedly select actions from finite sets, with payoffs determined by a fixed but unknown game matrix.
Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix blum2007external . In this paper, we focus specifically on a TPZSG setting with bandit feedback and propose two algorithms, ETC-TPZSG and ETC-TPZSG-AE, designed to minimize regret by learning pure strategies for both players through balancing of exploration and exploitation.
Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback. The variant of UCB algorithm for adversarial games under bandit feedback have been analyzed using worst-case regret bounds in o2021matrix . In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game. Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees. More recently, instance-dependent regret bounds have been investigated for the Tsallis-INF algorithm in TPZSGs ito2025instance . However an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
@@
This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs. While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to adversarial game settings, which makes this the first work to provide its detailed analysis in such a setting. A key motivation for studying ETC is its algorithmic simplicity allowing for a clear structure.
In particular, we propose two algorithms that adapt the ETC approach to TPZSG setting and present their instance-dependent regret analyses. The first one, ETC-TPZSG, follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure strategy, achieves an upper bound of $\tilde{O}(\Delta+\sqrt{T})$ The second algorithm, ETC-TPZSG-AE, incorporates an adaptive elimination approach, similar to auer2010ucb , by sequentially eliminating actions which, with high probability, are not part of an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action spaces of the players. Consequently, despite the simplicity of ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings. In both cases, the regret bounds depend on a suitable $\Delta$ notion, which varies depending on the type of the regret we are analyzing.
@@
This paper is organized as follows. Section 2 reviews related work. Section 3 introduces preliminaries and formalizes the two-player zero-sum game (TPZSG) setting. In Section 4, we present the ETC in a TPZSG setting, ETC-TPZSG, and analyze its regret. Section 5 introduces an elimination based algorithm, ETC-TPZSG-AE, that leverages the $\varepsilon$-NE property, and provides its regret analysis. Section 6 offers empirical results validating their theoretical performances. Finally, Section 7 discusses our conclusions and future research directions.
@@ 2 Related Work
The multi-armed bandit (MAB) problem, has been introduced by thompson1933likelihood and originally described by robbins1952some , has been widely explored and has become a crucial framework for online decision-making under uncertainty, as explored in works such as audibert2009exploration ; auer2002finite ; auer1995gambling ; auer2010ucb ; blum2007learning ; bubeck2012regret . The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff, has formalized by lai1985asymptotically and further explored in studies such as audibert2007tuning ; maes2012learning .
@@
A central challenge is identifying the action with the highest expected reward while minimizing suboptimal choices, requiring efficient exploration and statistical confidence. Several approaches guarantee low error probabilities within finite trials, aiming to minimize regret audibert2010best ; gabillon2012best . Since the player has no prior knowledge of expected rewards, a dedicated exploration phase is essential, has motivated numerous studies such as bubeck2009pure ; neu2015explore .
@@
The Explore-Then-Commit (ETC) algorithm, which has been introduced by perchet2016batched , is one of the fundamental approaches in the MAB setting. It explores each action a fixed number of times, then commits to the best-performing action for the remaining rounds. Due to its simplicity and effectiveness, variants of the ETC algorithm are widely used in various decision-making scenarios garivier2016explore ; yekkehkhany2019etc .
@@
Zero-sum games, are a fundamental concept in game theory. They describe competitive scenarios where one player’s gain exactly equals the other’s loss. First highlighted by v1928theorie ; vonneumanandmorgenstern , their theory and applications have been extensively explored blum2007learning ; cesa2006prediction ; washburn2014two . The Minimax Theorem in v1928theorie ; vonneumanandmorgenstern guarantees equilibrium existence in TPZSGs, but finding equilibria becomes significantly harder when the payoff matrix is unknown daskalakis2011near ; daskalakis2009complexity ; mazumdar2019finding ; zinkevich2007new .
A more general concept are Nash Equilibria nash1950equilibrium ; nash1951non . While these are identical to the minimax equilibria in TZPSGs, they are not as easy to compute in general games daskalakis2009complexity . The concept of $\varepsilon$-Nash Equilibrium provides an approximate solution where each player’s strategy is within an $\varepsilon$ tolerance level. daskalakis2009note used the $\varepsilon$-NE concept to efficiently approximate equilibrium in two-player games.
@@
In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, making it essential to estimate the unknown payoffs by previous observations. Various approaches, such as UCB o2021matrix and Tsallis-INF ito2025instance , have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret. Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments. o2021matrix focuses on worst-case regret analysis, while ito2025instance provides an instance-dependent regret analysis which achieves an upper bound of $O(\frac{\log T}{\Delta})$ in the case of a pure strategy equilibrium. Additionally, maiti2023instance investigates instance-dependent sample complexity bounds for zero-sum games under bandit feedback while cai2023uncoupled examines convergence rates to NE.
@@ 3 Preliminaries
In this section, we define the basic concept of a TPZSG such as payoff matrix, the regret definitions we focus on and the gaps $\Delta$ that we use for the analysis.
@@
Notation.
Throughout this paper, $O(\cdot)$ denotes an upper bound up to constant factors.
@@ 3.1 Two-Player Zero-Sum Games
$S_{x}$ and $S_{y}$ are the action sets of the row player and column player, respectively. Thus, the action space of the game is $S_{A}=S_{x}\times S_{y}$, so that the action pairs $(i,j)\in S_{A}$ where $i\in S_{x},j\in S_{y}$. We also define $m=|S_{x}|$ as the number of actions for the row player, and $l=|S_{y}|$ as the number of actions for the column player, while finally $N=ml=|S_{A}|$ is the total number of action pairs. The game itself is defined through a game matrix $A$, so that if the players choose the pair $(i,j)$ then the row player obtains a payoff of $A(i,j)$ and the column player $-A(i,j)$.
@@
Definition 3.1.
A pair of mixed strategies $(p^{*},q^{*})$ is a Nash equilibrium (NE) if for all strategies $p$ and $q$, it satisfies
@@
where $V^{*}={p^{*}}^{\top}Aq^{*}$ is the value of the game nash1950equilibrium . It means that $p^{*}$ and $q^{*}$ are optimal strategies for row and column players, respectively.
There always exists an optimal mixed strategy for both players that guarantees the best possible outcome by the Minimax Theorem v1928theorie ; vonneumanandmorgenstern . When a mixed strategy assigns a probability of one to a single action and zero to all others, it is referred to as a pure strategy. In our setting, we focus on a TPZSG where both players adopt pure strategies.
Definition 3.2 (Pure Nash Equilibria).
A pair $(i^{*},j^{*})$ is a pure NE if:
$i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)\quad\text{ and }\quad j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j),$
@@ 3.2 Games with Bandit Feedback
The game proceeds in rounds. At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$. The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, with the maximizing player obtaining $r_{t}$ and the minimizing player $-r_{t}$. We assume that both players observe each other’s action.
Since the algorithms do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played. We express the action pair played in round $t$ by $(i^{t},j^{t})$. After round $t$, we estimate each entry of the estimated game matrix by computing the average payoff for each action pair $(i,j)$ as
$\hat{A}_{t}(i,j)=\frac{1}{n_{ij,t}}\sum_{s=1}^{t}\mathbb{I}\{(i^{s},j^{s})=(i,j)\}r_{s}$ (3)
where $n_{ij,t}$ is the number of times the action pair $(i,j)$ played until round $t$, $r_{s}$’s are independently and identically distributed (iid) $\sigma$-subgaussian payoffs. We assume that the estimator is unbiased, meaning that $\mathbb{E}[\hat{A}]=A$.
@@ 3.3 Gap Notions
We note that $\Delta_{ij}^{\text{max}}\geq 0$ and $\Delta_{ij}^{\text{min}}\geq 0$, thus $\Delta_{ij}\geq 0$ for any action pair $(i,j)$. However, $\Delta_{ij}^{*}$ might be negative.
@@
$A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)\implies A(i^{*},j^{*})-A(i,j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)-A(i,j)$ (8)
$A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})\implies A(i^{*},j^{*})-A(i,j)\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})-A(i,j)$ (9)
Then, the inequality (8) implies that $\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$ and similarly from (9), we have $\Delta_{ij}^{*}\geq-\Delta_{ij}^{\text{min}}$. Therefore, we have $-\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$. On the other hand, since $\Delta_{ij}^{\text{min}}\geq 0$, we can write
@@ 3.4 Regret Notions
To characterize the performance of our algorithms, we consider several regret notions. Let begin with the following formulations, which are the external regret of the max and min player respectively:
$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\Big{[}\sum_{t=1}^{T}A(i,j^{t})-A(i^{t},j^{t})\Big{]}$
$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big{[}\sum_{t=1}^{T}A(i^{t},j^{t})-A(i^{t},j)\Big{]}$
where $i^{t}$ and $j^{t}$ are the actions selected by the row and column player in round $t$, respectively. $R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret from not choosing the column player’s best action for a given action $i$.
@@
$R_{T}^{\text{max}}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}],$ $R_{T}^{\text{min}}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}],$ (11)
@@
Definition 3.3 (External regret).
@@
External regret quantifies how much worse the players have performed compared to their best actions in hindsight. An alternative regret notion is the Nash regret, which measures the deviation of value from the Nash equilibrium:
Definition 3.4 (Nash regret).
@@
4 ETC-TPZSG Algorithm and Regret Analysis
In this section, we extend ETC algorithm to a two-person zero-sum game setting as ETC-TPZSG. Each player selects action from their corresponding finite action sets, which are known. The game is repeated over given time horizon $T$, and the true payoff matrix $A$ is unknown to both players. Instead, they receive observations of the payoffs based on their chosen actions, which are iid subgaussian random variables.
Let $k$ represent the number of times to explore each action pair $(i,j)\in S_{A}$. During the exploration phase, each player samples actions to estimate the expected payoffs, which enables us to construct an empirical payoff matrix using the observed payoffs. Then, in commit phase, they play according to the pure NE strategy derived from the estimated payoff matrix. After presenting the ETC-TPZSG algorithm, finally, we analyze its expected regret, measuring the performance due to limited information and suboptimal action pairs during the exploration phase.
As a result, the unknown payoff matrix is approximated by the ETC approach, as in Lattimore_Szepesvari_2020 for a bandit setting. The ETC algorithm consists of two phases: an exploration phase, where the player randomly samples actions to gather information with given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the gathered data. Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the committing phase the algorithm converges to an optimal action pair which refers to the pure NE.
Applying ETC in a zero-sum game setting presents some challenges compared to the standard MAB scenario. In bandit problems, after the exploration phase, a player simply commits to the arm with the highest estimated reward, which is typically effective if exploration is sufficient. However, in a zero-sum game with the pure NE, the objective is not to identify an action with high reward but to find a possibly true NE. Inadequate exploration can lead to poor estimates of payoffs, making it difficult to correctly identify such an equilibrium. In fact, due to inaccurate estimates, a pure NE might not even appear to exist, even when one does in the true game. As a result, the design of ETC approach in zero-sum game setting requires careful balancing of exploration phase.

### v2

@@ Abstract
We study a two-player zero-sum game in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback. We propose three algorithms based on the Explore-Then-Commit framework. The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration. Our objective is to demonstrate the applicability of ETC in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria. A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret of our proposed algorithms, which has received limited attention in the literature on zero-sum games. Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O(\log(T\Delta^{2})/\Delta)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap. Therefore, our results indicate that ETC-based algorithms perform effectively in adversarial game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
Keywords zero-sum games, regret analysis, bandit feedback
@@ 1 Introduction
This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in an adversarial setting with strictly opposing objectives, as first analyzed in [31, 32]. These games play a foundational role in game theory and online learning, capturing adversarial interactions in various applications such as economic competition and adversarial machine learning. In such settings, players repeatedly select actions from finite sets, with payoffs determined by a fixed but unknown game matrix.
Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix [7]. In this paper, we study a TPZSG setting with bandit feedback, where players observe each other’s actions, by learning pure strategies for both players through balancing of exploration and exploitation.
Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback. The variant of UCB algorithm for adversarial games under bandit feedback has been analyzed using worst-case regret bounds in [27]. In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game. Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees. More recently, instance-dependent regret bounds have been investigated for the Tsallis-INF algorithm in TPZSGs [18]. However, an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
@@
This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs. While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to adversarial game settings, which makes this the first work to provide its detailed analysis in such a setting. A key motivation for studying ETC is its algorithmic simplicity, which allows for a clear structure.
In particular, we propose three algorithms that adapt the ETC approach to the TPZSG setting and present their instance-dependent regret analyses. The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieves an upper bound of $O(\Delta+\sqrt{T})$ . The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players. Finally, the last algorithm extends the adaptive elimination approach through non-uniform exploration, which also achieves a regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Consequently, despite the simplicity of the ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings.
@@
Section 2 reviews related work. Section 3 introduces preliminaries and formalizes a two-player zero-sum game setting. In Section 4, we present the ETC in a TPZSG setting and analyze its regret. Section 5 introduces an adaptive elimination algorithm and provides its regret analysis. In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret bound. Section 7 offers empirical results to validate their theoretical performances. Finally, Section 8 discusses our conclusions and future research directions.
@@ 2 Related Work
The multi-armed bandit (MAB) problem, has been introduced by [30] and originally described by [29], has been widely explored and has become a crucial framework for online decision-making under uncertainty, as explored in works such as [3, 4, 5, 6, 8, 9]. The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff, has been formalized by [19] and further explored in studies such as [2, 21].
@@
A central challenge is identifying the action with the highest expected reward while minimizing suboptimal choices, which require efficient exploration and statistical confidence. Several approaches guarantee low error probabilities within finite trials, in order to minimize regret [1, 16]. Since the player has no prior knowledge of the expected rewards, a dedicated exploration phase is essential, which has motivated numerous studies such as [10, 26].
@@
The Explore-Then-Commit (ETC) algorithm, which has been introduced by [28], is one of the fundamental approaches in the MAB setting. It explores each action a fixed number of times, then commits to the best performing action for the remaining rounds. Due to its simplicity and effectiveness, variants of the ETC algorithm are widely used in various decision-making scenarios [17, 34].
@@
Zero-sum games represent a core concept in game theory. They describe competitive scenarios in which the gain of one player is exactly equal to the loss of the other. First highlighted by [31, 32], their theory and applications have been extensively explored [8, 12, 33]. The Minimax Theorem in [31, 32] guarantees the existence of equilibrium in TPZSGs, but finding equilibrium becomes significantly harder when the payoff matrix is unknown [13, 14, 23, 35].
Furthermore, the concept of Nash Equilibrium, which is introduced by [24, 25], describes optimal strategies of players in which no player is incentive to deviate. The study [14] shows that converging to such an equilibrium is challenging. The concept of $\varepsilon$-Nash Equilibrium provides an approximate solution where the strategy of each player is within a tolerance level $\varepsilon$. [15] uses this property to efficiently approximate equilibrium in two-player games.
@@
In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, which makes it essential to estimate the unknown payoffs by previous observations. Various approaches, such as UCB [27] and Tsallis-INF [18], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret. Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments. Additionally, [22] investigates instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
While [27] focuses on the worst-case regret analysis, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium. Compared to [18], our work introduces more refined notions of suboptimality gaps for game instances, whereas their regret analysis focuses solely on the gap between the NE and the action chosen by a player when the opponent plays the equilibrium action. Since their regret bound involves the square of the sum of suboptimality gaps, our instance-dependent analysis gives a tighter bound.
@@ 3 Preliminaries
In this section, we define the basic concept of a TPZSG such as payoff matrix, the regret definitions we focus on, and the gaps $\Delta$ that we use for the analysis.
@@ 3.1 Two-Player Zero-Sum Games
$S_{x}$ and $S_{y}$ are the action sets for the row player and column player, respectively. Thus, the action space of the game is $S_{A}=S_{x}\times S_{y}$, so that the action pairs $(i,j)\in S_{A}$ where $i\in S_{x},j\in S_{y}$. We also define $m=|S_{x}|$ as the number of actions for the row player and $l=|S_{y}|$ as the number of actions for the column player, while finally $N=ml=|S_{A}|$ is the total number of action pairs. The game itself is defined through a game matrix $A$, therefore if the players choose the pair $(i,j)$ then the row player obtains a payoff of $A(i,j)$ and the column player A(i,j)-A(i,j).
@@
Definition 1.
A pair of mixed strategies $(p^{*},q^{*})$ is a Nash Equilibrium if for all strategies $p$ and $q$, it satisfies
@@
where $V^{*}={p^{*}}^{\top}Aq^{*}$ is the value of the game [24]. This means that $p^{*}$ and $q^{*}$ are optimal strategies for row and column players, respectively.
There always exists an optimal mixed strategy for both players that guaranties the best possible outcome according to the Minimax Theorem [31, 32]. When a mixed strategy assigns a probability of one to a single action and zero to all others, it refers to a pure strategy. In our setting, we focus on a TPZSG where both players adopt pure strategies.
Definition 2 (Pure Nash Equilibria).
A pair $(i^{*},j^{*})$ is a unique pure NE if:
$i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)\quad\text{ and }\quad j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j),$ (2)
@@ 3.2 Games with Bandit Feedback
The game proceeds in rounds. At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$. The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, with the maximizing player obtaining $r_{t}$ and the minimizing player $r_{t}$. We assume that both players observe each other’s action.
Since the algorithms do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played. We express the action pair played in round $t$ by $(i_{t},j_{t})$. After round $t$, we estimate each entry in the estimated game matrix by computing the average payoff for each action pair $(i,j)$ as
$\hat{A}_{t}(i,j)=\frac{1}{n_{ij,t}}\sum_{s=1}^{t}\mathbb{I}\{(i_{s},j_{s})=(i,j)\}r_{s}$ (4)
where $n_{ij,t}$ is the number of times the action pair $(i,j)$ played until round $t$, $r_{s}$’s are independently and identically distributed (iid) $\sigma$-subgaussian payoffs. We assume that the estimator is unbiased, which means that $\mathbb{E}[\hat{A}]=A$.
@@ 3.3 Gap Notions
We note that $\Delta_{ij}^{\text{max}}\geq 0$ and $\Delta_{ij}^{\text{min}}\geq 0$, therefore $\Delta_{ij}\geq 0$ for any action pair $(i,j)$. However, $\Delta_{ij}^{*}$ might be negative.
@@
$A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)$
which implies
$A(i^{*},j^{*})-A(i,j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)-A(i,j).$ (9)
Moreover,
$A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})$
implies the following:
$A(i^{*},j^{*})-A(i,j)\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})-A(i,j).$ (10)
Then, the inequality (9) implies that $\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$ and similarly from (10), we have $\Delta_{ij}^{*}\geq-\Delta_{ij}^{\text{min}}$. Therefore, we have $\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$. On the other hand, since $\Delta_{ij}^{\text{min}}\geq 0$, we can write
@@ 3.4 Regret Notions
To characterize the performance of our algorithms, we consider several regret notions. Let begin with the following formulations, which are the external regret of the max and min players, respectively:
$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i,j_{t})-A(i_{t},j_{t})\Big]$
$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i_{t},j_{t})-A(i_{t},j)\Big]$
where $i^{t}$ and $j^{t}$ are the actions selected by the row and column player in round $t$, respectively. $R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of the row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret of not choosing the column player’s best action for a given action $i$.
@@
$R_{T}^{\text{max}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}],$
$R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}],$
@@
Definition 3 (External regret).
@@
External regret quantifies how much worse the players have performed compared to their best actions in hindsight. An alternative regret notion is Nash regret, which measures the deviation of value from the Nash equilibrium:
Definition 4 (Nash regret).
@@
4 ETC-TPZSG Algorithm
In this section, we extend the ETC algorithm to a two-person zero-sum game setting, called ETC-TPZSG. Each player selects action from their corresponding finite action sets, which are known. The game is repeated over a given time horizon $T$, and the true payoff matrix $A$ is unknown to both players. Instead, they receive observations of the payoffs based on their chosen actions, which are iid subgaussian random variables.
Let $k$ represent the number of times to explore each action pair $(i,j)\in S_{A}$. During the exploration phase, each player samples actions to estimate the expected payoffs, which enables us to construct an empirical payoff matrix using the observed payoffs. Then, in the commit phase, they play according to the pure NE strategy derived from the estimated payoff matrix. Finally, we analyze the expected regret of the algorithm, measuring the performance due to limited information and suboptimal action pairs during the exploration phase.
As a result, the unknown payoff matrix is approximated by the ETC approach, as in [20] for a bandit setting. The ETC algorithm consists of two phases: an exploration phase, in which the player randomly samples actions to gather information with a given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the observations. Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the committing phase the algorithm converges to an optimal action pair which refers to the pure NE.
Applying ETC in a zero-sum game setting presents some challenges compared to the standard MAB scenario. In bandit problems, after the exploration phase, a player simply commits to the arm with the highest estimated reward, which is typically effective if exploration is sufficient. However, in a zero-sum game with pure NE, the objective is not to identify an action with high reward but to find a possibly true NE. Inadequate exploration can lead to poor estimates of payoffs, making it difficult to correctly identify such an equilibrium. As a result, the design of the ETC approach in the zero-sum game setting requires careful balance of the exploration phase.